# Contextual AD Narration with Interleaved Multimodal Sequence

### v1

@@ B More Ablation Studies
B.1 Ablation on the structure of visual mapping network
Table 7: Ablation on the structure of visual mapping network. latent denotes the number of learnable vectors in our visual mapping network. Experiments are conducted with Uni-AD(GPT-2-B32).
Visual mapping network RLnormal-$\uparrow$ Cnormal-$\uparrow$ R@5/16normal-$\uparrow$
@@ B.1 Ablation on the structure of visual mapping network
As stated in Sec. 3.2 of the main paper, there are multiple reasons why we choose a multi-layer transformer encoder with a fixed number of learnable vectors as our mapping network. Here we compare our visual mapping network with two different visual mapping designs: MLP and multi-layer transformer encoder without learnable vectors. Results in  Tab. 7 show that no interaction between video frames(MLP as visual mapping network) gets the worst performance. Allowing interaction between video frames(transformer encoder as visual mapping network) brings better results, but the length of visual embeddings is limited to be consistent with the number of frames(8 in our experiments). Our visual mapping network with 30 learnable vectors performs the best.
B.2 Ablation on the impact of sharing visual mapping network
Table 8: Ablation on the impact of sharing visual mapping network. Share? shows whether we encode both video and image with one single visual mapping network.
Methods Share? RLnormal-$\uparrow$ Cnormal-$\uparrow$ R@5/16normal-$\uparrow$
@@ B.2 Ablation on the impact of sharing visual mapping network
Since in Uni-AD, the structure of video mapping network is the same as image mapping network, we in this section study the impact of using a single visual mapping network to encode both video and image. The results are shown in  Tab. 8. We can see that encoding video and image with two separate mapping networks is important to Uni-AD(GPT-2-B32), while not necessary for Uni-AD(LLaMA-L14). This reflects that when visual features and LLM are good enough, images and videos can be mixed together for training the mapping network.
Table 9: Ablation on more interaction between image and video. Inter-I.? shows whether we concatenate character’s image and the video clip as input of image mapping network. Inter-V.? shows whether we concatenate all character images and the video clip as input of video mapping network.
Methods Inter-I.? Inter-V.? RLnormal-$\uparrow$ Cnormal-$\uparrow$ R@5/16normal-$\uparrow$
@@
B.3 Ablation on more interaction between image and video
In the main paper, we encode video and image into visual tokens and apply the frozen LLM for interaction between video and image. To study whether more interaction between image and video can benefit AD generation, we replace the input of visual mapping network as concatenation of image and video. Specifically, we replace the input with concatenation of current character’s image and the video for image mapping network. For video mapping network, we replace the input with concatenation of all recognized characters and the video. In this way, we investigate whether the visual mapping network can extract better character and video representations by more interaction. The results are shown in  Tab. 9, which indicates that more interaction between image and video for visual mapping can not further benefit our Uni-AD.
C Additional Qualitative Analyses
@@
Figure 6 shows more qualitative results of Uni-AD on the MAD-eval dataset. Note that our character-refinement module can not only recognize AD-related characters, but also serve as a character information denoiser. For example, in sample (a) where characters Graham and Merrill do not appear in the video clip but are included in the character bank, our character-refinement module removes these noises and provides more precise character information. Though AD without character-refinement module also focuses on describing the female police officer, it can not figure out who she is since there are noises in the initial character bank, thus mistakes Caroline as Merrill’s mom. In sample (c), we find that with more learnable vectors, our model can take the female character who appears at the beginning into AD generation. However, the female character is just ignored by our Uni-AD with fewer learnable vectors.

### v2

@@ Appendix C More Ablation Studies
C.1 Ablation on visual mapping network
As stated in Sec. 3.2 of the main paper, there are multiple reasons why we choose a multi-layer transformer encoder with a fixed number of learnable vectors as our mapping network. Here we compare our visual mapping network with two different visual mapping designs: MLP and multi-layer transformer encoder without learnable vectors. Results in  Tab. S3 show that no interaction between video frames (MLP as visual mapping network) gets the worst performance. Allowing interaction between video frames(transformer encoder as visual mapping network) brings better results, but the length of visual embeddings is limited to be consistent with the number of frames(8 in our experiments). Our visual mapping network with 30 learnable vectors performs the best.
Table S3: Ablation on the structure of visual mapping network. latent denotes the number of learnable vectors in our visual mapping network. Experiments are conducted with Uni-AD(GPT-2-B32)
[…]
@@ C.1 Ablation on visual mapping network
Table S4: Ablation on the impact of sharing visual mapping network. Share? shows whether we encode both video and image with one single visual mapping network.
[…]
@@
C.2 Ablation on sharing visual mapping network
Since in Uni-AD, the structure of video mapping network is the same as image mapping network, we in this section study the impact of using a single visual mapping network to encode both video and image. The results are shown in  Tab. S4. We can see that encoding video and image with two separate mapping networks is important to Uni-AD(GPT-2-B32), while not necessary for Uni-AD(LLaMA-L14). This reflects that when visual features and LLM are good enough, images and videos can be mixed together for training the mapping network.
C.3 Ablation on image-video interaction
In the main paper, we encode video and image into visual tokens and apply the frozen LLM for interaction between video and image. To study whether more interaction between image and video can benefit AD generation, we replace the input of visual mapping network as concatenation of image and video. Specifically, we replace the input with concatenation of current character’s image and the video for image mapping network. For video mapping network, we replace the input with concatenation of all recognized characters and the video. In this way, we investigate whether the visual mapping network can extract better character and video representations by more interaction. The results are shown in  Tab. S5, which indicates that more interaction between image and video for visual mapping can not further benefit our Uni-AD.
Table S5: Ablation on more interaction between image and video. Inter-I.? shows whether we concatenate character’s image and the video clip as input of image mapping network. Inter-V.? shows whether we concatenate all character images and the video clip as input of video mapping network.
[…]
@@ C.3 Ablation on image-video interaction
C.4 Ablation on the Threshold in Character-Refinement Module
We conduct ablation study on the impact of threshold in Character-Refinement Module to our Uni-AD and the results in Tab. S6 show that the threshold has a considerable impact on AD generation. High threshold may lead to excessive loss of character information thus gets poor results.
Table S6: Ablation on the Threshold value in Character-Refinement Module.
@@ C.4 Ablation on the Threshold in Character-Refinement Module
0.3 16.5 25.7 52.4 0.5 16.8 27.3 53.3
0.7 16.2 24.8 55.4 0.9 12.8 16.1 54.3
Appendix D Additional Qualitative Analyses
Figure S1: Qualitative analysis on character-refinement module and contextual information. Movies are selected from (a): Signs (2002), (b): The Ides of March (2011).
[…]
Figure S1 and  Figure S2 shows more qualitative results of Uni-AD on the MAD-eval dataset. Note that our character-refinement module can not only recognize AD-related characters, but also serve as a character information denoiser. For example, in sample (a) where characters Graham and Merrill do not appear in the video clip but are included in the character bank, our character-refinement module removes these noises and provides more precise character information. Though AD without character-refinement module also focuses on describing the female police officer, it can not figure out who she is since there are noises in the initial character bank, thus mistakes Caroline as Merrill’s mom. In sample (c), we find that with more learnable vectors, our model can take the female character who appears at the beginning into AD generation. However, the female character is just ignored by our Uni-AD with fewer learnable vectors.